OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME SOLVE THIS PROBLEM. THE OWNER OF MARINA BOUGHT 6 CANOES TO BE USED AS RENTAL AT A COST OF $650 PER CANOE. THEY BORROWED 70% OF THE TOTAL COST FOR 8 MONTHS AT 10 1/2% INTEREST. FIND THE AMOUNT DUE USING SIMPLE INTEREST.

OpenStudy (anonymous):

ie. lol. okay so 6 x 650. hold up. someone replayed to my question.

OpenStudy (anonymous):

OKKK

OpenStudy (anonymous):

do you want the answer ot help?

OpenStudy (anonymous):

cause i iknow i like when people give me the answer

OpenStudy (anonymous):

6*650= 3900

OpenStudy (anonymous):

okay.. no find 70 0f that number x/3900 = 70/100 Cross mulpiply & divide

OpenStudy (anonymous):

sorry. meant to say now

OpenStudy (anonymous):

IM CONFUSED ON THIS PART

OpenStudy (anonymous):

what class are you in?

OpenStudy (anonymous):

ARE YOU DIVIDING

OpenStudy (anonymous):

Im In Algebra

OpenStudy (anonymous):



OpenStudy (anonymous):

Apply Simple Interest formula!

OpenStudy (anonymous):

truu. that too

OpenStudy (anonymous):

no i dont know the formula

OpenStudy (anonymous):

70 *3900/100=27300

OpenStudy (anonymous):

Your job is looking up the Simple Interest formula, post it here, then I show you how to apply :)

OpenStudy (anonymous):

are you talking about the I = P*R*T

OpenStudy (anonymous):

3900 PRINCIPLE Rate 0.70 Time 8/12

OpenStudy (anonymous):

so, 2730*.105* 8/12

OpenStudy (anonymous):

so 8/12 can be reduce too 2/3

OpenStudy (anonymous):

I dont have a caluclator at this moment.

OpenStudy (anonymous):

Use the one from your PC!

OpenStudy (anonymous):

i just calculated 2730*.105=286.65

OpenStudy (anonymous):

how would i do 8/12

OpenStudy (anonymous):

i cant do 2/3 on this calculator

OpenStudy (anonymous):

Multiply by 2, then divide by 3 ..

OpenStudy (anonymous):

191.1

OpenStudy (anonymous):

how can you round it off too nearest cents

OpenStudy (anonymous):

would it be 191.10

OpenStudy (anonymous):

Amount due = P + I

OpenStudy (anonymous):

191+2730

OpenStudy (anonymous):

=2921.1

OpenStudy (anonymous):

so 2921.1 is the answer

OpenStudy (anonymous):

The concept is Simple Interest, and it indeed fairly simple!

OpenStudy (anonymous):

ok